# EvAlignUX: Advancing UX Research through LLM-Supported Exploration of Evaluation Metrics

### v1

@@ 3.1. System Design
Based on the design goals, EvAlignUX , finalizes with three main panels, as shown in Figure 1. To illustrate interactions in EvAlignUX , we present a walk-through with Sarah, a UX researcher designing an AI counseling chatbot for alcohol addiction. The user journey is likened to a ”shopping” process, where Sarah selects ideal evaluation plans for her proposed system.
User Journey:
Once landing on EvAlignUX , Sarah inputs her project details. She names the project “Mindful Drinking”, selects “brainstorming” and “designing the study” as the current project statuses, and provides a project description, evaluation plans,
and details the expected outcomes
After entering all details, she clicks “Generate” and proceeds to the main interaction page.
@@ 3.1.1. Project Ideation Panel (DG1)
This panel extracts relevant indexes to categorize the work within the existing literature. In the rapidly evolving field of UX research, evaluators often face challenges in understanding the context and relevance of newly proposed systems (Ledo et al., 2018). To enhance the UX evaluator’s understanding of the context in which a system operates and its significance within the broader field of UX research, we built on categorizations from prior literature reviews and key guiding literature for UX of AI design practices (Zheng et al., 2022; Liao et al., 2020b; Vermeeren et al., 2010; Kieffer et al., 2019). We leverage these established frameworks to integrate critical factors into the system output. This output is designed to automatically map the proposed system, based on the provided narrative, into categories including paradigms, application domains, modalities, system features, design novelty, design methods, human-AI relationship types, stakeholders, social scales, and theoretical frameworks.
User Journey:
Sarah views a list of indexed recommendations from EvAlignUX , noting that some are selected while others are not. She realizes ”Addiction Recovery” isn’t checked under ”Application Domain,” so she selects it (A). She also checks ”Family Members” as the evaluation target stakeholder, which she had previously overlooked. After making these changes, she clicks ”Regenerate” (B) followed by ”Compare Metrics” (C). This process updates her a list, showing added, retained, and removed metrics (D), sparking her curiosity about the new metrics.
@@ 3.1.2. Metric Explorer Panel (DG1, DG2)
This panel recommends relevant evaluation metrics and related literature while enabling users to explore additional metrics through both list-view and graph-view interactions, helping them formulate a comprehensive evaluation plan. Prior literature has highlighted that users are often excluded from the process of defining evaluation criteria for a system, with most participants consulted only at the end of the design process to assess the system’s effects, rather than being actively involved in setting these criteria. This exclusion contradicts the goal of democratizing the design process (Bossen et al., 2016; Zheng and Huang, 2023). Inspired by this challenge, we propose a metric recommendation feature that introduces UX evaluation metrics into design materials as a design or evaluation input. Given that graph-based interaction is a commonly used and effective approach for recommendation systems (Ali et al., 2020), we propose incorporating a graph-view alongside the traditional list-view. This graph-view will allow users to visualize UX metrics and their correlating metrics, fostering inspiration and enabling more informed filtering and selection of metrics. Furthermore, prior literature outlines various methods for collecting metrics, such as system logs, surveys, and interviews (Zheng et al., 2022). To enhance user understanding, we will provide context and examples of how these metrics have been derived and applied.
User Journey:
Sarah starts by reviewing suggested metrics in the list (A) and is intrigued by ”goal attainment,” defined as assessing users’ perceptions of achieving their recovery goals through the system (D). She clicks the metric info button (C) to explore how this metric is measured—through surveys, system logs (E), and interviews—and reviews related papers on each method’s application and findings (F). Interested, she adds it to her shopping cart (B) and uses the Metric Graph feature to explore additional metrics.
@@ 3.1.3. Outcomes and Risks Panel (DG2, DG3)
Assists users in framing their research contributions based on previous studies and identifies potential pitfalls by referencing databases of prior AI risks and incidents. Technologies have positively transformed society, but they have also led to undesirable consequences not anticipated at the time of design or development. Previous HCI research explored using insights into past undesirable consequences in the world to help researchers and practitioners gain awareness anticipate potential adverse effects and reflect on their own experiences with technology (Pang et al., 2024). Inspired by this, we propose a panel for generating related research outcomes and risks based on existing research and real-world cases. We sourced case studies from a database indexing real AI incidents documented by journalists — the AI, Algorithmic, and Automation Incident and Controversy (AIAAIC) repository that previous papers have used (Lee et al., 2024b). This approach aims to strengthen UX evaluation metrics by incorporating research outcomes based on literature, and relevant risks grounded in real-world incidents.
As Sarah adds and removes metrics from her shopping cart, she notices that the results in the Outcome feature update dynamically. She selects outcomes from prior studies that catch her interest. Then, she explores potential risks associated with her proposed UX evaluation plan by clicking on the Risk feature. “All set!”, said Sarah. She then proceeds with exporting the data and conducts a final review of her generated UX evaluation plan and UX outcomes report.
@@ 3.2. System Backend and Implementation
In this section, we describe the process of preparing the data repository, system backend and implemenation, as well as the techniques used for prompting and making recommendations. The frontend of the system is implemented using ReactJS111https://github.com/facebook/react, TypeScript and CSS. The backend uses Python with FastAPI222https://github.com/tiangolo/fastapi/ as the RESTful API server framework. The data is represented as a Knowledge Graph using Neo4j333https://github.com/neo4j Graph Database to model the complex relations among research papers (Verma et al., 2023; Kejriwal, 2022; Wang et al., 2018). We use the GPT-4o model as our LLM engine and set the temperature to 0.7 for all components. The embeddings in the system are generated using text-embedding-3-small, with FAISS444https://github.com/facebookresearch/faiss as the vector database for data retrieval. Additionally, Langchain555https://github.com/langchain-ai/langchain is used for prompting the GPT-4o model.
@@ 3.2.1. Data Preparation
In this section, we present how the scientific database was prepared. Based on several UX evaluation survey papers, e.g.,(Zheng et al., 2022; Inan Nur et al., 2021; Vermeeren et al., 2010), we established an initial repository of evaluation metrics. Two authors expanded this repository by incorporating additional metrics identified in the reviewed papers. Definitions for all metrics, including the newly added ones, were either directly sourced from these surveys or developed collaboratively by the authors to ensure completeness and coherence.
@@
After building the repository of evaluation metrics, we created a paper database from the same survey papers and utilized GPT-4o (OpenAI et al., 2024) to build up more fine-grained information associated with the evaluation metrics. We prompted GPT with each full text of the papers and asked GPT to provide a narrative of the system developed in the papers, along with the challenges or the research gaps that the papers addressed. We provided index definitions and asked GPT to generate a list of ten descriptive indices based on the content of the papers. These descriptive indices include: interaction paradigms, application domain, modality, system features, design novelty, design methods, human-AI relationship types, stakeholders involved, social scale and theoretical frameworks used. The definitions for these indices can be found in Appendix A.1.
@@
Between the metrics and the papers in the database, we also created the usage information of each metric in relevant paper. For example, a metric usage follows the following format: ”This paper uses the [X] metric to evaluate users’ [Y] towards [Z technology]. It was found that [S findings related to Y].” [X] is the metric identified in 1, [Y] is the aspect of user interaction being evaluated, [Z] is the technology or system being evaluated, and [S] is the findings related to the user interaction being evaluated.
@@
We then crawled metadata and appended metadata to each paper. The metadata includes the uid of the paper, the title, authors, the abstract, the date of publication, uids of papers that cited this paper, keywords, the publication venue, the publisher and the affiliation of authors. Besides, we also leverage the existing AI Incident Database as a repository for risks (McGregor, 2020) and extracted relevant incidents and risks. EvAlignUX ’s backend recommends UX metrics based on an initial System Description. After a series of user interactions, it generates a new Evaluation Plan and UX Outcome using LLM and data from research papers via a Knowledge Graph, representing complex entity relations derived from source documents.
@@ 3.2.2. Index Generation and derivation of Knowledge Graph
A list of indexes is generated by EvAlignUX using the initial system description provided by the user. The indexes are intended to standardize the system description into fixed categories. Indexes are generated by providing the System Description, examples of other generated indexes, and corresponding definitions of each index as input to the LLM. Subsequently, the user is recommended three additional index values corresponding to each category. These indexes and the initial system description are used to retrieve a list of metrics from our Knowledge Graph.
The Knowledge Graph serves as the foundation for recommending metrics, built using the Neo4j Graph Database. We structured our data as a graph to capture the intricate relationships within our dataset (Verma et al., 2023; Kejriwal, 2022; Wang et al., 2018), enabling efficient information extraction and retrieval of global contextual information (Edge et al., 2024). The backing dataset of research papers is organized into distinct clusters of communities (Fortunato, 2010) based on their index similarity, citations, and common metrics using the Louvain Community detection algorithm (De Meo et al., 2011). The graph consists of three different types of nodes, each characterized by distinct properties. Table 5 in the Appendix provides a detailed overview of the distinct nodes and their associated properties. The three unique nodes are as follows:
@@
Outcome Node: This node is attached to the Metric node via an edge and contains the outcome and other information corresponding to each metric used in a paper. Since the metric node could be shared between multiple paper nodes, this node stores the unique information of a metric corresponding to a research paper.
All nodes are unique and do not contain duplication. Furthermore, weighted edges exist between the nodes. For instance, there exist three edges amongst each pair of paper nodes, namely,
[SHARED_METRIC], [CITES], and [CITED_BY]. The
[SHARED_METRIC] edge exists between two paper nodes if they share at least one common metric. Similarly, if any paper A cites another paper B, there exists a [CITES] edge from paper node A to paper node B, and a [CITED_BY] edge from paper node B to paper node A. Edges of type [CITES] and [CITED_BY] are assigned relatively higher weight, compared to [SHARED_METRIC] edge, by multiplying the edge weight with a constant value, due to stronger relation amongst the papers because of the citation relation, based on an increase in the modularity score as discussed in (Blondel et al., 2008). The weight assigned to these edges is influenced by the relative importance of each index category in evaluating the similarity between research papers. Index categories that indicate greater similarity receive higher weights, which are then multiplied by the similarity score between index values of the categories for each paper, contributing to the overall edge weight. Additionally, each Paper node is connected to a Metric node via a non-weighted edge of type [HAS_METRIC], which is then connected to an Outcome node via another non-weighted edge of type [HAS_OUTCOME].
@@ 3.2.3. Metrics Recommendation
The provided system description and its corresponding generated indexes extract metrics from the Knowledge Graph. The system description and indexes are vectorized, and the most similar paper is determined using cosine similarity in the existing Knowledge Graph. The community for the most similar paper is determined from the above Knowledge Graph and consequently, all metrics linked to the paper nodes belonging to the above community are extracted. These metrics, along with the user system description and generated indexes are used to further filter metrics, by prompting a LLM. Additionally, each metric is further classified into three metric collection methods, based on their usage in previous research papers, using the Knowledge Graph. These metrics are displayed in a list view and a graphical view to the user.
@@ 3.2.4. Potential risks and metric outcomes
Each metric in the Knowledge Graph is associated with outcomes corresponding to the research paper. Subsequently, these results are retrieved from the Knowledge Graph based on the final metrics list, and displayed to the user to filter, based on their relevance to the user’s system description. Each outcome is specific to a research paper and is mapped to all the metrics associated with the research paper. Consequently, these outcomes are further considered while generating the Potential Evaluation Plan and the Potential UX Outcome.
In addition to the recommended metric outcomes, our system generates potential risks associated with the user system, using the AIAAIC repository. We extract the top three incidents that have a Euclidean distance of less than 0.5 by performing a vector-based similarity search, using FAISS, between the user system description, and the AI system description present in the AI Incident dataset. Subsequently, we extract risks associated with these AI systems, and provide them along with the user system description to a LLM, to filter the most relevant risks for the system. The risks identified are contextualized based on the initial system description, with corresponding URLs for related incidents also available to the user.
Once the metrics have been finalized by the user, along with the corresponding metric outcomes, a Potential Evaluation Plan and Potential UX Outcome is recommended to the user. The Potential Evaluation Plan includes detailed instructions, regarding the usage of each metric and the methods for measuring each metric, based on the user’s System Description and initial evaluation plan. The Potential Evaluation Plan is generated by providing the system description, a list of selected metrics, the specified metric outcome, and methods used to previously evaluate the UX Metrics, to a LLM, and subsequently, prompting the LLM to generate the Potential Evaluation Plan. The Potential UX Outcomes are generated by considering the list of selected metrics, desired results from the metric outcomes, initial system description, and also the risks associated with the system.

### v2

@@ 3.1. System Design
Based on the DG2, EvAlignUX comprises with three main panels, as shown in Figure 1. To illustrate interactions in EvAlignUX , we present a walk-through with Sarah, a UX researcher designing an AI counseling chatbot for alcohol addiction. The user journey is likened to a ”shopping” process, where Sarah selects ideal evaluation plans for her proposed system. See Figure 2 for User Journey 1.
User Journey 1:
Upon landing on EvAlignUX , Sarah inputs her project details. She names the project “Mindful Drinking”, selects “brainstorming” and “designing the study” as the current project statuses, and provides a project description, evaluation plans, and details the expected outcomes. After entering all details, she clicks “Generate” and proceeds to the main interaction page.
The graph shows a snapshot of the EvalignUX system landing page (left) and a user journey (right). On the left, the landing page displays a form for creating a new project. This includes fields like “Project Name”, “Project Status”, “System Description”, “System UX Evaluation”, “UX Evaluation Result”. At the bottom, a “Generate” button allows the user to proceed to the next step. On the right, the accompanying text outlines the user journey. It explains that upon landing on EvalignUX, the user (Sarah) inputs project details, names the project “Mindful Drinking”, selects “brainstorming” and “designing the study” as the current project statuses, provides a system description, UX evaluation plan, and expected outcomes. After entering these details, she clicks “Generate” to continue to the main interaction page.
@@ 3.1.1. Project Ideation Panel (DG1)
This panel extracts relevant indexes to categorize the work within the existing literature. In the rapidly evolving field of UX research, evaluators often face challenges in understanding the context and relevance of newly proposed systems (Ledo et al., 2018). To enhance the UX evaluator’s understanding of the context in which a system operates and its significance within the broader field of UX research, we built on categorizations from prior literature reviews and key guiding literature for UX in AI design practices (Zheng et al., 2022; Liao et al., 2020b; Vermeeren et al., 2010; Kieffer et al., 2019). We leverage these established frameworks to integrate critical factors into the system output. This output is designed to automatically map the proposed system based on the provided narrative, into categories including paradigms, application domains, modalities, system features, design novelty, design methods, human-AI relationship types, stakeholders, social scales, and theoretical frameworks. See Appendix 10 for details. See Figure 3 for User Journey 2.
User Journey 2:
Sarah views a list of indexed recommendations from EvAlignUX , noting that some are selected while others are not. She realizes that ”Addiction Recovery” isn’t checked under ”Application Domain,” so she selects it (A). She also checks ”Family Members” as the evaluation target stakeholder, which she had previously overlooked. After making these changes, she clicks ”Regenerate” (B) followed by ”Compare Metrics” (C). This process updates her list, showing added, retained, and removed metrics (D), sparking her curiosity about the new metrics.
The graph shows a snapshot of the EvalignUX Project IDEATION Panel and a user journey (right). On the left, the IDEATION Panel displays indexed recommendations under indexes such as “Application Domain”, with options like “Addiction Recovery” and “Mental Health”. Users can select or deselect relevant indexes, for example, checking “Addiction Recovery” (marked as ’A’) to refine the system evaluation. Below, buttons for “Regenerate Metrics”’ and “Compare Metrics” allow the user to update and compare evaluation metrics. The Metrics comparison window (center) shows the changes of metrics recommended by the system before and after users change the indexes. On the right, the text outlines Sarah’s journey as she reviews indexed recommendations from EvalignUX. Noticing that “Addiction Recovery” is unchecked under “Application Domain”, she selects it (A). She also selects “Family Members” as an evaluation target, which she had previously missed. After making these changes, Sarah clicks “Regenerate” (B) and “Compare Metrics” (C), which generates a list of newly added, retained, or removed metrics (D), piquing her curiosity about the updated evaluation metrics.
@@ 3.1.2. Metric Explorer Panel (DG1, DG2)
This panel recommends relevant evaluation metrics and related literature while enabling users to explore additional metrics through both a list-view and graph-view interactions, helping them formulate a comprehensive evaluation plan. See Figure 4 for User Journey 3.
Prior literature has highlighted that users are often excluded from the process of defining evaluation criteria for a system, with most participants consulted only at the end of the design process to assess the system’s effects, rather than being actively involved in setting these criteria. This exclusion contradicts the goal of democratizing the design process, which seeks to empower all stakeholders to contribute meaningfully throughout the design life-cycle (Bossen et al., 2016; Zheng and Huang, 2023). Inspired by this challenge, we propose a metric recommendation feature that integrates UX evaluation metrics into design materials as inputs for design or evaluation. Given that graph-based interaction is a commonly used and effective approach for recommendation systems (Ali et al., 2020), we propose incorporating a graph-view alongside the traditional list-view. This graph-view will allow users to visualize UX metrics and their related metrics, fostering inspiration and enabling more informed filtering and selection of metrics. Furthermore, prior literature outlines various methods for collecting metrics, such as system logs, surveys, and interviews (Zheng et al., 2022). To enhance user understanding, we will provide context and examples of how these metrics have been derived and applied.
User Journey 3:
Sarah starts by reviewing the suggested metrics in the list (A) and is intrigued by the ”goal attainment”, defined as assessing users’ perceptions of achieving their recovery goals through the system (D). She clicks the metric info button (C) to explore how this metric is measured—through surveys, system logs (E), and interviews—and reviews related papers on each method’s application and findings (F). Interested, she adds it to her shopping cart (B) and uses the Metric Graph feature to explore additional metrics.
The graph shows a snapshot of the Metrics Explorer Panel (left) and a user journey (right). The metric list (A) allows users to browse through various metrics, and by clicking the metric info button (C), users can access detailed information about each metric, including definitions (D), collection methods (E), and additional research papers or findings (F). Metrics can also be added to a “shopping cart” (B) for later use. On the right, the user journey describes Sarah’s experience. She begins by reviewing the suggested metrics in the list (A) and is particularly interested in “goal attainment”, which is defined as assessing users’ perceptions of achieving their recovery goals through the system (D). She clicks the metric info button (C) to explore how this metric is measured, discovering data collection methods like surveys, system logs (E), and interviews. Sarah then reviews related research papers that provide insights into each method’s application and findings (F). After gaining an understanding, she adds the metric to her shopping cart (B) and uses the Metric Graph feature to explore additional metrics.
@@ 3.1.3. Outcomes and Risks Panel (DG2, DG3)
Assists users in framing their research contributions based on previous studies and identifying potential pitfalls by referencing databases of prior AI risks and incidents. Technologies have positively transformed society, but they have also led to undesirable consequences not anticipated at the time of design or development. Previous HCI research has explored using insights from past undesirable consequences to help researchers and practitioners gain awareness of potential adverse effects and reflect on their own experiences with technology (Pang et al., 2024). Inspired by this, we propose a panel for generating related research outcomes and risks based on existing research and real-world cases. We sourced case studies from a database indexing real AI incidents documented by journalists — the AI, Algorithmic, and Automation Incident and Controversy (AIAAIC) repository, which previous papers have used (Lee et al., 2024b). This approach aims to strengthen UX evaluation metrics by incorporating literature-based research outcomes and relevant risks grounded in real-world incidents.
User Journey 4: As Sarah adds and removes metrics from her shopping cart, she notices that the results in the Outcome feature update dynamically. She selects outcomes from prior studies that catch her interest. Then, she explores potential risks associated with her proposed UX evaluation plan by clicking on the Risk feature. “All set!” said Sarah. She then exports the data and conducts a final review of her generated UX evaluation plan and UX outcomes report.
@@ 3.2. System Backend and Implementation
In this section, we describe the process of preparing the data repository, system backend, and implementation, as well as the techniques used for prompting and making recommendations. The frontend of the system is implemented using ReactJS111https://github.com/facebook/react, TypeScript and CSS. The backend uses Python with FastAPI222https://github.com/tiangolo/fastapi/ as the RESTful API server framework. The data is represented as a Knowledge Graph using the Neo4j333https://github.com/neo4j Graph Database to model the complex relations among research papers (Verma et al., 2023; Kejriwal, 2022; Wang et al., 2018). We use the GPT-4o model as our LLM engine and set the temperature to 0.7 for all components. The embeddings in the system are generated using text-embedding-3-small, with FAISS444https://github.com/facebookresearch/faiss as the vector database for data retrieval. Additionally, Langchain555https://github.com/langchain-ai/langchain is used for prompting the GPT-4o model.
@@ 3.2.1. Data Preparation
In this section, we present how the scientific database was prepared. First, we established an initial repository of evaluation metrics based on the definitions specified in prior literature (e.g., (Zheng et al., 2022)). Soon, we recognized that the literature review focused on conversational AI interactions; thus two authors expanded the UX evaluation metrics to include interactions in other modalities by incorporating additional metrics that were widely applied beyond conversational interactions (Inan Nur et al., 2021; Vermeeren et al., 2010). For example, we expanded the repository by incorporating metrics related to ethics and risks, such as ”communication fairness,” ”user control,” and ”privacy concerns,” which address broader dimensions of responsible interaction design. We do not claim that the repository is complete or exhaustive. It will continue to grow as new HCI literature and UX evaluation metrics emerge. Definitions for all metrics, including the newly added ones, were either directly sourced from these surveys or developed collaboratively by the authors to ensure completeness and coherence.
@@
After building the repository of evaluation metrics, we created a paper database from the same survey papers and utilized GPT-4o (OpenAI et al., 2024) to generate more fine-grained information associated with the evaluation metrics. We prompted GPT with the full text of each paper and asked it to provide a narrative of the system described in the papers, along with the challenges or the research gaps that the papers addressed. We provided index definitions and asked GPT to generate a list of ten descriptive indices based on the content of the papers. These descriptive indices include: interaction paradigms, application domain, modality, system features, design novelty, design methods, human-AI relationship types, stakeholders involved, social scale, and theoretical frameworks used. The definitions for these indices can be found in Appendix A.1.
@@
Between the metrics and the papers in the database, we also created usage information for each metric in the relevant papers. For example, a metric usage is formatted as follows: ”This paper uses the [X] metric to evaluate users’ [Y] towards [Z technology]. It was found that [S findings related to Y].” [X] represents the metric identified in the paper, [Y] is the aspect of user interaction being evaluated, [Z] is the technology or system being evaluated, and [S] is the findings related to the user interaction being evaluated.
@@
We then crawled metadata and appended it to each paper. The metadata includes the paper’s UID, title, authors, abstract, publication date, UIDs of papers that cited it, keywords, publication venue, publisher, and authors’ affiliations. Besides, we also leverage the existing AI Incident Database as a repository for risks (McGregor, 2020) and extracted relevant incidents and risks. EvAlignUX ’s backend recommends UX metrics based on an initial System Description. After a series of user interactions, it generates a new Evaluation Plan and UX Outcome using an LLM and data from research papers via a Knowledge Graph, which represents complex entity relations derived from source documents.
This figure illustrates the workflow of EvalignUX, a system that supports HCI researchers in developing UX evaluation plans by integrating LLM support, structured data, and user inputs. Users provide a project description, initial evaluation plan, and initial outcome, which the system processes to generate relevant indexes and UX metrics using a Knowledge Graph. These metrics help define potential UX outcomes, while risk considerations are assessed using data from risk samples and an AI Risk repository. Finally, the system synthesizes the selected metrics, risks, and user inputs to generate a structured UX evaluation plan and outcome, enabling researchers to explore diverse evaluation approaches while considering impact and potential risks.
@@ 3.2.2. Index Generation and derivation of Knowledge Graph
Our system, EvAlignUX , generates a list of indexes using the initial system description provided by the user. The indexes are intended to standardize the system description into fixed categories. Indexes are generated by providing the System Description, examples of other generated indexes, and the corresponding definitions of each index as input to the LLM. Subsequently, the system recommends three additional index values for each category. These indexes and the initial system description are used to retrieve a list of metrics from our Knowledge Graph.
The Knowledge Graph, built using the Neo4j Graph Database, serves as the foundation for recommending metrics. We structured our data as a graph to capture the intricate relationships within our dataset (Verma et al., 2023; Kejriwal, 2022; Wang et al., 2018), enabling efficient information extraction and retrieval of global contextual information (Edge et al., 2024). The backing dataset of research papers is organized into distinct clusters of communities (Fortunato, 2010) based on their index similarity, citations, and common metrics, using the Louvain Community detection algorithm (De Meo et al., 2011). The graph consists of three different types of nodes, each characterized by distinct properties. Table A.2 in the Appendix provides a detailed overview of the distinct nodes and their associated properties. The three unique nodes are as follows:
@@
Outcome Node: It is attached to the Paper Node and the Outcome Node via several edges. The node is shared between two Paper Nodes if they measure the same metric. Since the metric node could be shared between multiple paper nodes, this node stores the unique information of a metric corresponding to each metric as used in a research paper.
All nodes are unique and do not duplicate. Furthermore, weighted edges exist between nodes. For instance, there exist three edges between each pair of paper nodes, namely, [SHARED_METRIC], [CITES], and [CITED_BY]. The [SHARED_METRIC] edge exists between two paper nodes if they share at least one common metric. Similarly, if any paper A cites another paper B, there exists a [CITES] edge from paper node A to paper node B, and a [CITED_BY] edge from paper node B to paper node A. Edges of type [CITES] and [CITED_BY] are assigned relatively higher weight, compared to [SHARED_METRIC] edge. This is achieved by multiplying the edge weight with a constant value, due to stronger relation among the papers because of the citation relation, based on an increase in the modularity score as discussed in (Blondel et al., 2008). The weight assigned to these edges is influenced by the relative importance of each index category in evaluating the similarity between research papers. Index categories that indicate greater similarity receive higher weights. These weights are then multiplied by the similarity score between index values of the categories for each paper, contributing to the overall edge weight. Additionally, each Paper node is connected to a Metric node via a non-weighted edge of type [HAS_METRIC], which is then connected to an Outcome node via another non-weighted edge of type [HAS_OUTCOME].
@@ 3.2.3. Metrics Recommendation
The provided system description and its corresponding generated indexes are used to extract metrics from the Knowledge Graph. The system description and indexes are vectorized, and the most similar paper is determined using cosine similarity on the existing Knowledge Graph. The community for the most similar paper is determined from the aforementioned Knowledge Graph and consequently, all metrics linked to the paper nodes belonging to that community are extracted. These metrics, along with the user system description and generated indexes, are used to further filter metrics, by prompting an LLM. Additionally, each metric is further classified into three metric collection methods, based on its usage in previous research papers, using the Knowledge Graph. These metrics are displayed in a list view and a graphical view to the user.
@@ 3.2.4. Potential risks and metric outcomes
Each metric in the Knowledge Graph is associated with outcomes corresponding to the research paper. Subsequently, these results are retrieved from the Knowledge Graph based on the final metrics list and displayed to the user so that they can filter them based on relevance to the user’s system description. Each outcome is specific to a research paper and is mapped to all metrics associated with that paper. Consequently, these outcomes are further considered when generating the Potential Evaluation Plan and the Potential UX Outcome.
In addition to the recommended metric outcomes, our system generates potential risks associated with the user system using the AIAAIC repository. We extract the top three incidents that have a Euclidean distance of less than 0.5 by performing a vector-based similarity search, using FAISS between the user system description and the AI system description present in the AI Incident dataset. Subsequently, we extract risks associated with these AI systems, and provide them along with the user system description to a LLM to filter the most relevant risks for the system. The identified risks are contextualized based on the initial system description, and corresponding URLs for related incidents are also provided to the user.
Once the metrics have been finalized by the user, along with the corresponding metric outcomes, a Potential Evaluation Plan and Potential UX Outcome are recommended to the user. The Potential Evaluation Plan includes detailed instructions on the usage of each metric and the methods for measuring them based on the user’s system description and initial evaluation plan. The potential evaluation plan is generated by providing the user’s system description, a list of selected metrics, the specified metric outcomes, and the methods previously used to evaluate the UX metrics to an LLM, and then prompting the LLM to generate the plan. The Potential UX Outcomes are generated by considering the list of selected metrics, the desired results from the metric outcomes, the initial System Description, and the risks associated with the system.